OpenStudy (anonymous):

Simplify cos2 Θ(1 + tan2 Θ).

OpenStudy (anonymous):

cos2 Θ + sin2 Θ cos2 Θ 1 −1

OpenStudy (anonymous):

@ganeshie8

OpenStudy (solomonzelman):

\(\large\color{blue}{ \cos^2θ(1+\tan^2θ) }\) \(\large\color{blue}{ \cos^2θ(1+\frac{\sin^2θ}{\cos^2θ}) }\) \(\large\color{blue}{ \cos^2θ+\sin^2θ }\) \(\large\color{blue}{ 1}\)

OpenStudy (anonymous):

thankyou @SolomonZelman

OpenStudy (solomonzelman):

anytime